as Bank of Frank Assets Liabilities Total $25,000 Deposits $150,000 Reserves: Excess Required Reserves Loans $100,000 Securities $25,000 The Required Reserve Ratio is 10% Select all that apply:

media%2F7cd%2F7cd7d758-0d44-4895-b655-b7
Answer #1

Required reserve = 10% of total deposits = 10% of 150,000 = 15000

Excess reserve = Actual reserve - Required reserve = 25000 - 15000 = 10,000

Bank can provide loan of 150,000 - 15,000 = 135,000

Money supply = Excess reserve x Multiplier = 10,000 x 1/Reserve requirement = 10,000 x 1/0.10 = 100,000

a) This bank is required to hold $ 15000 in reserves.

b) The money supply would increase if the Federal Open Market Committee purchased this bank's bonds.

Purchase of government bonds increases money supply. Commercial banks influence money supply by controlling amount of loans in the economy.